OpenStudy (anonymous):

Solve for x: 4x – 2 = x + 4 A. x = 1/2 B. x = -1/2 C. x = 2 D. x = -2

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

C

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

4X - 2 = X +4 Add 2 to both sides so you can get 4x = x+6 Keep x on one side by subtracting x from both sides 3x = 6 Divide by 3 to get x alone and you get x = 2